# REDUCIO! Generating 10241024 Video within 16 Seconds using Extremely Compressed Motion Latents

### v2

@@ 5 Conclusion
Video generation has shown promising results with many potential applications, but is still trapped by the unaffordable computation cost. In this paper, we explored how to effectively cut the overhead by reducing the latent size. Particularly, we found the video can be encoded into extremely compressed latents with the help of a content image prior, where the latent codes only need to represent the motion variables. With this observation, we designed the Reducio-VAE to compress videos to smaller latents. Using this powerful Reducio-VAE, we trained the Reducio-DiT that enables fast high-resolution video generation. Our method is also compatible with other accelerating techniques, e.g. rectified flow, allowing for further speedup, where we leave it for future exploration.
Limitations and future work. While Reducio-DiT shows a strong capacity of being consistent with the center frame, the generated videos are still relatively short (16 frames of 16 fps, i.e., 1 second) for real-world applications. Given the length of the video, the magnitude of the motion is also somehow limited in generated videos. We believe in the potential of adapting our works for longer video generation and will explore it as the future work.
Reducio is purely a research project. Currently, we have no plans to incorporate Reducio into a product or expand access to the public. We will also put Microsoft AI principles into practice when further developing the models. In our research paper, we account for the ethical concerns associated with text-image-to-video research. To mitigate issues associated with training data, we have implemented a rigorous filtering process to purge our training data of inappropriate content, such as explicit imagery and offensive language, to minimize the likelihood of generating inappropriate content.
@@ Appendix A Implementation details
Reducio-VAE. We demonstrate the detailed architecture of Reducio in Fig. 8. Generally, we follow the training strategies of SD-VAE [40]. We employ a tailored version of PatchGAN [18] based on 3D convolutions and optimize the model with loss, KL loss, perceptual loss [68], and GAN loss. While initializing the 2D encoder and the 3D VAE with SD-VAE pre-trained weights accelerates convergence, we find that freezing the 2D encoder leads to worse performance than training the full parameters. We follow LaMD [16] to feed the motion latent into a normalization layer to obtain the output of the VAE encoder. Therefore, in the stage of diffusion training, we use a scale factor of 1.0, which is multiplied by the input latent as inputs into DiT.

### v3

@@ 5 Conclusion
Video generation has shown promising results with many potential applications while it is still trapped by the unaffordable computation cost. In this paper, we explored how to effectively cut the overhead by reducing the latent size. In particular, we found that video can be encoded into extremely compressed latents with the help of a content image prior, where the latent codes only need to represent the motion variables. With this observation, we designed the Reducio-VAE to compress videos to smaller latents. Using this powerful Reducio-VAE, we trained the Reducio-DiT that enables fast high-resolution video generation. Our method is also compatible with other accelerating techniques, e.g., efficient attention, rectified flow, diffusion distillation, and etc., allowing for further speedup, where we leave it for future exploration.
Limitations and future work. While Reducio-DiT maintains strong consistency with the center frame, the number of training videos is still relatively limited for good instruction following performance. Also, 16-frame training samples somehow limit the magnitude of the motion in generated videos. As Reducio is much more compute-friendly with increased video frames and extended training data. Therefore, we believe in the potential of adapting our work for longer video generation on large-scale training and will explore it as future work.
Acknowledgement This work was supported in part by the National Natural Science Foundation of China (Grant 62032006 and Grant 62472098).
@@ Appendix A Implementation details
Reducio-VAE. We demonstrate the detailed architecture of Reducio in Fig. 10. Generally, we follow the training strategies of SD-VAE [45]. We employ a customized version of PatchGAN [21] based on 3D convolutions and optimize the model with loss, KL loss, perceptual loss [82], and GAN loss. While initializing the 2D encoder and the 3D VAE with SD-VAE pre-trained weights accelerates convergence, we find that freezing the 2D encoder leads to worse performance than training the full parameters. We follow LaMD [19] to feed the motion latent into a normalization layer to obtain the output of the VAE encoder. Therefore, in the stage of diffusion training, we use a scale factor of 1.0, which is multiplied by the input latent as input into DiT.
@@
Appendix B More ablations
| Conv | PSNR | SSIM | |
| 64 | 2d | 30.22 | 0.86 |
| 64 | 3d | 35.51 | 0.94 |
Using 3d VAE in Reducio-VAE helps to reconstruct videos in a better quality. We keep to 1 and to 32 and implement VAE with 2d convolutions. During decoding, we duplicate the middle frame condition for times to fuse with the latent of each frame respectively. As shown in Tab. 10, Reducio-VAE with 3d convolution outperforms its counterpart with 2d convolution. We believe that 3d convolution facilitates the VAE to model consistent motion and capture spatiotemporal differences.
| Content Frame | PSNR | SSIM |
| n/a | 27.91 | 0.80 |
| random | 31.72 | 0.87 |
| middle | 35.88 | 0.94 |
Using middle frame in Reducio-VAE. The content frame in Reducio-VAE provides a strong content prior and hence leads to a promising reconstruction performance. On the other hand, relying on any given frame as the content image may not generalize perfectly in all scenarios, especially when certain entities appear only briefly or outside the chosen frame. As shown in Tab. 11, we choose the middle frame by default as it serves as a more stable and robust content guidance due to its temporal centrality. Meanwhile, Reducio-VAE without condition achieves significantly worse results in both PSNR (-7.97) and SSIM (-0.14). In consequence, the Reducio-DiT framework without condition-based 3D VAE leads to unsatisfactory results featured with blurry frames and obvious visual defects.
| Model | latent shape | PSNR | SSIM | |
| DC-AE [7] | 32 | 30.68 | 0.70 | |
| Reducio-VAE | 16 | 35.56 | 0.97 |
Comparison between DC-AE and Reducio-VAE. We compare Reducio-VAE with DC-AE [7] on the Pexel test split with resolution of . As shown in the Table below, Reducio-VAE outperforms DC-AE on PSNR and SSIM by 4.88 and 0.27, respectively, highlighting the advantage of our framework in video domain.
| Attn | FVD | IS |
| 2d + 1d | 382.2 | 32.4 |
| 3d | 337.6 | 34.1 |
@@ Appendix B More ablations
A dog wearing a Superhero outfit with red cape flying through the sky.
A dog wearing a Superhero outfit with red cape flying through the sky.
Using joint spatiotemporal 3D attention in Reducio-DiT outweighs using factorized spatial and temporal attention (i.e., 2D + 1D attention) in generation quality. Interestingly, we observe that factorized attention leads to a faster convergence of training loss. However, with the same training step, as shown in Tab. 13, factorized attention lags behind its counterpart with joint 3D attention for 45 in FVD. We suppose the possible reason is that 2D + 1D scheme demands adding additional temporal layers and performs factorized self-attention on a small set of tokens e ach, making it hard to model smooth open-set motion with the light computation. In contrast, 3D attention directly exploits the original parameters and collaborates all spatiotemporal tokens.
Quantitative results. We display the detailed performance comparison on VBench [20] in Tab. 15 and Tab. 14. Despite using only 3.2K A100 GPU hours and 5.4M training samples, Reducio-DiT achieves a promising semantic score of 78.00, beating a range of state-of-the-art LDMs. While the most recent models such as WAN [58] and CausVid [75] achieve higher overall scores than Reducio-DiT, we argue that our model uses a much smaller scale of training data and has a relatively small model scale, i.e., 1.2B.